Analyze the IR spectra of bromobenzene
Benzene, bromo- INFRARED SPECTRUM TRANSMITTANCE 1000 3000 2000 Wavenumber (cm-1)
Answer #1

Analysis of IR Spectra of bromobenzene

3100cm-1 - C-H(aromatic) stretching vibrations

* Weak combinations and overtones bands are appears in the range of 2050 - 1650cm-1 .

1600-1500cm-1 ​​​​ - C=C stretching of aromatic ring

1070-1000cm-1 - =C-H bending absorption

750-650cm-1  - C-Br stretching absorption bands

please rate my answer thankyou.